# Enabling Advanced Land Cover Analytics: An Integrated Data Extraction Pipeline for Predictive Modeling with the Dynamic World Dataset

### v1

@@ II-A Land Use Land Cover Datasets
Many works in urbanization forecasting do not leverage the existing LULC labels available in open source datasets, such as the Dynamic World dataset, and instead train their own models to first predict the LULC labels  [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. The use of our pipeline in conjunction with open source LULC datasets should eliminate the need for custom LULC labelling.
@@ II-B Land Use Land Cover Dataset Processing Pipeline
Despite the difficulty and steep learning curve associated with working with satellite imagery, to the extent of our knowledge, there does not exist another open source pipeline like ours which includes extraction, pre-processing, and representation of LULC data. As we will discuss in Section II-C, many works prefer to use their own LULC labeling rather than using datasets like Dynamic World, and therefore rely on custom downstream pre-processing implementations. There is a package, GeeMap, which offers tools for satellite data analysis and visualization. As we will discuss, we have build upon and extended this package as its initial implementation does not include the end to end pipeline offered in our work.
@@ III-B1 Data Extraction and pre-processing
$d=2r\arcsin\left(\sqrt{\sin^{2}\left(\frac{\Delta\varphi}{2}\right)+\cos(\varphi_{1})\cos(\varphi_{2})\sin^{2}\left(\frac{\Delta\lambda}{2}\right)}\right)$ (1)
@@
The ImageExporter module is responsible for using the Earth Engine Python API and builds upon the Python GeeMap package to implement the data extraction process. Rather than using the native JavaScript approach to interact with Google Earth Engine Code Editor, we devised a scalable and development-ready data extraction pipeline using Python. This module adeptly converts Python code into Javascript queries, dispatched to the Earth Engine API, facilitating the extraction of images that are subsequently stored in a designated Google Drive folder.
@@
This process entails extracting multiple images for the same year, resulting in an image collection corresponding to multiple timestamps of one fishnet tile, and then applying pixel-level aggregation operations. Some of the available aggregation operations are average, minimum, maximum, median, etc. The advantage of retrieving an aggregated image over a single image lies in noise reduction, with noise stemming from various factors such as cloud cover or sensor noise. The Google Earth Engine offering has built-in functions for aggregating image collections into a single image, however, we opted to incorporate image aggregation in our ImageCorrector module to avoid needing to re-download images in case a different aggregation technique becomes necessary.
@@
Figure 4: Dynamic World exports: (a) before the correction and (b) after imputing missing pixels.
@@ IV Application Case Study and Results
Depending on the training complexity and performance needed, one can opt to train on either tabular data or images directly. Our work accommodates both modeling approaches, and the impressive performance of both XGBoost and XGCLM models highlights the quality and stability of our pre-processing pipeline.
@@ V Interesting Research Avenues
Our current pipeline is primarily structured to handle LULC satellite images. However, a potential enhancement to this study would be to expand the pipeline’s capabilities to seamlessly process various other types of satellite imagery, each encoding different information. In this case, the output would contain multiple customizable information layers, providing a more comprehensive and tailored analysis.
@@ Acknowledgments
This paper was prepared for informational purposes by the Artificial Intelligence Research group of J.P. Morgan Chase & Co. and its affiliates (“JP Morgan”), and is not a product of the Research Department of JP Morgan. JP Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein.

### v2

@@ II-A Land Use Land Cover Datasets
Many works in urbanization forecasting do not leverage the existing LULC labels available in open source datasets, such as the Dynamic World dataset, and instead train their own models to first predict the LULC labels  [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. The use of our pipeline in conjunction with open source LULC datasets should eliminate the need for custom LULC labelling. It should be noted that LULC labels extracted from such datasets, while of high quality, are not ground truth since they are generated by deep learning classification algorithms. As a consequence, we do not dismiss alternative approaches that may provide higher accuracy or robustness. Rather, our use of open source datasets reflects a practical concern: many teams lack the resources (computational or time) to generate large-scale, representative LULC labels, making datasets like Dynamic World a valuable baseline for global-scale applications. We also note that there may be other open source LULC datasets which outperform Dynamic World in some regions and time periods. Our pipeline could be easily amended to include these datasets, but we leave this line of work for future studies.
@@ II-B Land Use Land Cover Dataset Processing Pipeline
Despite the difficulty and steep learning curve associated with working with satellite imagery, to the extent of our knowledge, there does not exist another open source pipeline like ours which includes extraction, pre-processing, and representation of LULC data. Many works prefer to use their own LULC labeling rather than using datasets like Dynamic World, and therefore rely on custom downstream pre-processing implementations.
For instance, some methods  [10] rely on georeferencing imagery and image-to-image registration to align data into a common coordinate system, followed by the use of commercial software (e.g., eCognition Developer) for object-based image analysis to derive LULC classifications. In contrast, our pipeline eliminates the need for these additional steps by directly extracting images in a consistent coordinate system. Additionally, our approach is entirely open-source, built in Python, and does not depend on proprietary systems.
Other methods employ supervised classification approaches, such as using historical imagery and manually digitized polygons as training data for classifiers  [11, 20]. These approaches rely on the assumption that classifiers trained on smaller, localized datasets will generalize effectively to broader, unseen regions—a limitation our pipeline avoids. On the other hand, unsupervised methods, like the segmentation-based approach in  [22], utilize Convolutional Neural Networks (CNNs) for feature clustering but often require manual corrections to remove residual noise. This manual intervention significantly limits scalability, especially when applied to large geographic regions.
There is a package, GeeMap, which offers tools for satellite data analysis and visualization. As we will discuss, we have build upon and extended this package as its initial implementation does not include the end to end pipeline offered in our work.
@@ III-B1 Data Extraction and pre-processing
Altough our pipeline was developed using images extracted from the Dynamic World dataset, it is not heavily dependent on it and could easily be amended to use a different source or include additional datasets.
@@
The fishnet approach does not compete with the existing Google Earth Engine (GEE) capabilities, but rather complements them. While GEE provides robust capabilities for selecting and processing specific regions, the fishnet approach addresses key challenges related to scalability, standardization, and ease of integration into machine learning workflows.
  • Scalability: GEE’s region-based processing works well for small areas, but faces challenges with large regions spanning multiple administrative units such as counties or zip codes due to their irregular size. The fishnet divides the area into uniform grid cells, enabling consistent granularity and seamless processing. Even using the Rectangle geometry within GEE does not work well; large areas cannot be easily exported, often necessitating manual segmentation. Although GEE’s built-in methods may be effective for extracting data from a few counties, significant modifications would need to be made to allow for extraction of data at scale. This is where the fishnet excels.
  • Standardization: Machine learning models require input data to be consistent in size and format; the fishnet ensures that all extracted image patches are of equal dimensions, eliminating the need for additional resizing or cropping steps that would be necessary using administrative boundaries like counties or zip codes. Even using GEE’s Rectangle geometry can result in non-uniform sizes, as the image dimensions depend on the absolute latitude of a region. For example, a square region of 0.1 degrees by 0.1 degrees in latitude and longitude would have different dimensions when exported as an image if it is near one of the poles versus near the equator. This irregularity is handled by the fishnet. There is a method within GEE that allows for tiled exporting of large areas, but generates web-friendly map tiles optimized for visualization, rather than data to be ingested by a machine learning model, unlike the fishnet we have proposed.
  • Ease of Processing: The fishnet grid allows for systematic and parallelized extraction of data across the entire study region. This design facilitates streamlined integration with machine learning pipelines, where uniform inputs are critical for efficient training and inference. To the extent of our knowledge, there does not exist a built-in GEE function which offers extraction of very large scale regions into uniformly sized areas with a mapping between the extracted data and the tile location.
@@
$\small d=2r\arcsin\left(\sqrt{\sin^{2}\left(\frac{\Delta\varphi}{2}\right)+\cos(\varphi_{1})\cos(\varphi_{2})\sin^{2}\left(\frac{\Delta\lambda}{2}\right)}\right)$ (1)
@@
The ImageExporter module is responsible for using the Earth Engine Python API and builds upon the Python GeeMap package to implement the data extraction process. Rather than using the native JavaScript approach to interact with Google Earth Engine Code Editor, we devised a scalable and development-ready data extraction pipeline using Python. This module adeptly converts Python code into Javascript queries, dispatched to the Earth Engine API, facilitating the extraction of images that are subsequently stored in a designated Google Drive folder. By combining GEE’s robust data extraction capabilities with the fishnet’s systematic segmentation, this approach enhances scalability, standardization, and usability for large-scale, machine-learning-oriented spatial analyses.
@@
This process entails extracting multiple images for the same year, resulting in an image collection corresponding to multiple timestamps of one fishnet tile, and then applying pixel-level aggregation operations. Some of the available aggregation operations are average, minimum, maximum, median, etc. The advantage of retrieving an aggregated image over a single image lies in noise reduction, with noise stemming from various factors such as cloud cover or sensor noise.
The assumption is that aggregation will lead to better signal due to its ability to smooth out short-term variations or anomalies. However, in highly dynamic environments, such as during rapid land-use changes or natural disasters, aggregation may inadvertently mask critical signals. To address this, our method can be adapted to account for such dynamics by employing several techniques, including:
  • Selective Aggregation: Aggregating over a subset of high-quality, temporally relevant images.
  • Change Detection: Identifying temporal outliers or sudden deviations to emphasize areas experiencing rapid change.
  • Dynamic Granularity: Adjusting the aggregation window based on the expected rate of change in the region.
If downstream machine learning models are used, additional strategies can be considered:
  • External Factors as Features: External factors can be incorporated as additional features or channels in the deep learning architecture, facilitating model learning.
  • Avoiding Aggregation: A more granular time horizon can be used, feeding the model many more images and avoiding aggregation altogether.
Each of these approaches has trade-offs, and it is crucial to consider the specific use case when deciding which method to apply.
The Google Earth Engine offering has built-in functions for aggregating image collections into a single image, however, we opted to incorporate image aggregation in our ImageCorrector module to avoid needing to re-download images in case a different aggregation technique becomes necessary.
@@
Figure 4: Dynamic World exports: left before the correction; right after imputing missing pixels.
@@ IV Application Case Study and Results
To address the evaluation of our pipeline with traditional approaches, we implemented a Cellular Automata (CA) model, which is widely used in the literature for spatio-temporal forecasting in remote sensing. Cellular Automata models consist of a grid of cells that evolve through discrete time steps according to simple, local rules based on the states of neighboring cells, leading to complex global patterns. This CA model was trained using the output of our preprocessing pipeline. Due to its pixel-by-pixel nature, the CA model is computationally intensive and was trained on a limited dataset, resulting in modest performance metrics (R²: 0.42). However, it is important to note that computations for CA can be parallelized, which would result in significant speed-ups. In our case, we did not implement such optimizations as the primary focus of this paper was to demonstrate the flexibility of our preprocessing pipeline rather than optimizing the CA approach. While these results do not match the performance of the XGBoost and XGCLM models, they demonstrate the pipeline’s adaptability to traditional approaches like CA.
This highlights the pipeline’s flexibility, enabling the extraction of data from various modalities and supporting both traditional and advanced modeling techniques. Depending on the training complexity and performance needed, one can opt to train on either tabular data or images directly. Our work accommodates many different modeling approaches, and the impressive performance of both XGBoost and XGCLM models highlights the quality and stability of our pre-processing pipeline.
@@ V Interesting Research Avenues
Our current pipeline is primarily structured to handle Dynamic World satellite images. In future studies, we hope to include other LULC datasets. Another potential enhancement to this study would be to expand the pipeline’s capabilities to seamlessly process various other types of satellite imagery, each encoding different information. In this case, the output would contain multiple customizable information layers, providing a more comprehensive and tailored analysis.
@@
Additionally, we acknowledge the role of external factors such as new regulations, city plans, or unforeseen socio-economic developments in shaping land cover dynamics. While our work focuses on historical data as a proof-of-concept case study, the proposed framework is inherently flexible and extensible, allowing for the inclusion of exogenous features that can represent these external factors. For example, the neural network architecture can accommodate additional information as input features or channels. Data such as proximity to major roads, distance to city centers, land use restrictions (e.g., private property, protected areas, or reservoirs), or planned infrastructure developments can be encoded into the model. These features could be incorporated as auxiliary inputs or directly encoded into the spatial grid as additional channels. For instance, new urban development regulations might be represented as categorical features at the tile level, indicating development eligibility or restrictions. Similarly, planned city infrastructure, such as roads or parks, could be encoded with specific values to guide the model’s predictions. Since it is complex to manually obtain and encode such information for large regions, large language models (LLMs) could be leveraged to analyze regulatory documents, city planning reports, and socio-economic updates to determine which geographic regions may be impacted. These insights could then be encoded as features for the forecasting model, enabling it to dynamically adapt to evolving conditions.
@@ Acknowledgments
This paper was prepared for informational purposes in part by the Artificial Intelligence Research group of JPMorgan Chase & Co. and its affiliates (”JP Morgan”) and is not a product of the Research Department of JP Morgan. JP Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.